OpenStudy (anonymous):

*** I will give you a medal if you help/answer *** The speed of the current in a river is 6 mph. A ferry operator who works that part of the river is looking to buy a new boat for his business. Every day, his route takes him 22.5 miles against the current and back to his dock, and he needs to make this trip in a total of 9 hours. He has a boat in mind, but he can only test it on a lake where there is no current. How fast must the boat go on the lake in order for it to serve the ferry operator’s needs?

OpenStudy (anonymous):

Take 22.5 miles divide by 9 hrs to find speed on water without current and add 6mph to the speed ?

OpenStudy (anonymous):

is this for sure how you do it.. ??

OpenStudy (anonymous):

i got 8.5

OpenStudy (anonymous):

thats how I would do it. think of it this way, u add 6 mph to the boat and it cancels out the river current, so your remaining speed is the speed needed to cross the river 22.5 miles in 9 hrs

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let x = speed of the boat in still water now let's find the distance equation for the boat when it goes against the current d = rt d = (x - 6)*t ... the speed gets slowed down against the current, so that's why it drops to x-6 22.5 = (x-6)*t now solve for t 22.5/(x-6) = t t = 22.5/(x-6) So it takes 22.5/(x-6) hours to go 22.5 miles upstream ------------------------------------------- Now let's find the equation when the boat goes with the current d = rt d = (x + 6)*t the speed gets faster with the current, so that's why it bumps up to x+6 22.5 = (x + 6)*t 22.5/(x + 6) = t t = 22.5/(x + 6) So it takes 22.5/(x+6) hours to go 22.5 miles downstream -------------------------------------------------------- You want the two times to add to 9 hours, so... 22.5/(x-6) + 22.5/(x + 6) = 9 solve this for x to get your answer

OpenStudy (anonymous):

how do i solve for x ?

OpenStudy (anonymous):

find a LCM to cancel out the equations ? or what ?

jimthompson5910 (jim_thompson5910):

good, multiply everything by the LCD to clear out the fractions

jimthompson5910 (jim_thompson5910):

the LCD is (x-6)(x+6)

OpenStudy (anonymous):

so i would multiply everything by 135 ?

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

no (x-6)(x+6)

OpenStudy (anonymous):

im confused ?

OpenStudy (anonymous):

so i take 22.5(x-6)(x+6)

jimthompson5910 (jim_thompson5910):

start with the first fraction